OpenStudy (anonymous):

plss answer this Use derivatives of limits y=7-4x/5+2x

OpenStudy (anonymous):

you have to find y'?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Use "derivative of a fraction", so, your y=f/g, then f=7-4x and g=5+2x. Then y'=(f'g-fg')/g^2

OpenStudy (anonymous):

Use quotient rule

OpenStudy (anonymous):

its not chain rule method .. its derivatives of limits

OpenStudy (anonymous):

(-70-4x)/25

OpenStudy (anonymous):

how did u get that ?? plss show us solution so i can understand it

OpenStudy (anonymous):

you sure destiny? For me it's -34/(5+2x)^2 D:

OpenStudy (anonymous):

My answer is wrong, I used a quotinent rule

OpenStudy (anonymous):

\[\frac{ 7-4(x+Deltax) }{ 5+2(x+Deltax) }\] this should be the simplified.. im confused

OpenStudy (anonymous):

y=(7-4x)/(5+2x) or y=7-4x/5+2x ???

OpenStudy (anonymous):

\[\frac{ 7-4x }{ 5+2x } the problem :\]

OpenStudy (anonymous):

okay, that's different, it will start with \[\lim \frac{ \frac{7-4(x+h)} {5+2(x+h)} - \frac {7-4x} {5+2x} }{ h } h\to 0\]

OpenStudy (anonymous):

h=delta X

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Your result should be the same. The difference is that you're obtaining the derivative by the limit's definition of it.

OpenStudy (anonymous):

why dont you use quotient rule much more easier

OpenStudy (anonymous):

is chain rule and derivatives of limits is the same ??

OpenStudy (anonymous):

it is not chain rule it is quotient rule lol ya it is the same and faster

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

All the derivative rules were obtained from the definition, so, yes, if you can use 'em, use 'em xD

OpenStudy (anonymous):

unless you are required to use the classic way lol

OpenStudy (anonymous):

tnx for all the help